An automobile on an expressway is accelerating at 25 ft/s2. If it started from rest and has been accelerating for
5.56 s., how far has it travelled during this time of acceleration?
1
Expert's answer
2020-05-11T20:12:04-0400

The distance traveled by a car

"d=v_0t+\frac{at^2}{2}\\
=0\times 5.56+\frac{25\times 5.56^2}{2}=386 \:\rm ft"